. If there are 5,400 students in University ABC, 1,425 in the College of Nursing, 865 in the College of Medicine, 1,250 in College of Management, and 1,860 in the College of Education how many will be the samples using the stratified sampling design?


1
Expert's answer
2021-09-30T02:57:32-0400

Let us assume that the student to be asked for a survey be 500

College of Nursing "\frac{1425}{5400}*500=132" students


College of Medicine "\frac{865}{5400}*500=80" students


College of Management "\frac{1250}{5400}*500=116" students


College of Education "\frac{1860}{5400}*500=172" students